OpenStudy (anonymous):

my question is: Integral ( 1 / (1 - sec x) )

OpenStudy (amistre64):

what if you add zero to the top? like say (sec-sec) ? then split the fraction

OpenStudy (anonymous):

Umm make it easy for me,, if you can give me the full answer will be better.

OpenStudy (amistre64):

\[\int \frac{1}{1-sec}\] \[\int \frac{1+sec-sec}{1-sec}\] \[\int \frac{1-sec}{1-sec}+\frac{sec}{1-sec}\] \[\int 1+\frac{sec}{1-sec}\] or we might be able to say sec(x) = u to begin with and construct an equivalent integral from us

OpenStudy (amistre64):

or if we multiply by the conjugate that might get it into something doable

OpenStudy (amistre64):

\[\int \frac{1}{1-sec}\] \[\int \frac{1}{1-sec}*\frac{1+sec}{1+sec}\] \[\int \frac{1+sec}{1+sec^2}\] but maybe not lol

OpenStudy (anonymous):

my teacher use this way: put u = some thing... then complate

OpenStudy (amistre64):

u = tan(x/2) perhaps?

OpenStudy (amistre64):

i never do well with us and trigs combined ... gives me a headache

OpenStudy (anonymous):

\[\frac{\cos{x}}{\cos{x} - 1}\] \[ \frac{\cos ^{2}\frac{x}{2} - \sin^{2}\frac{x}{2}}{\cos ^{2}\frac{x}{2} - \sin^{2}\frac{x}{2} -1}\]

OpenStudy (anonymous):

I have got a headache from this question

OpenStudy (amistre64):

i tried to follow wolframs steps and got even more discombobulated

OpenStudy (anonymous):

\[\frac{1 - \tan^{2} \frac{x}{2}}{1 - \tan^{2}\frac{x}{2} - \sec^{2}\frac{x}{2}}\]

OpenStudy (anonymous):

ishaan94 there is no cos in the question!! give me more explan.

OpenStudy (anonymous):

\[\tan \frac{x}{2} = u\] \[ \frac12 \sec^{2}\frac{x}{2}= \frac{du}{dx}\] \[dx = \frac{2\times du}{\sec^{2}\frac{x}{2}}\] \[\frac{1 - u^2}{\left(1 - u^2 -(1 + u^2)\right)\left(1 + u^2\right)}2 du\] \[\frac{1 - u^2}{-2u^2(1 + u^2)}2du\] \[-\frac{1-u^2}{u^2(1 + u^2)}du\]

OpenStudy (anonymous):

Use partial fraction or something else...I am outta here Good Luck

OpenStudy (anonymous):

OK thanks.. I will try.